# Doubly-charmed pentaquark states in a mass splitting model

### v1

@@ III.4 Spectra and decay widths of $ccss\bar{n}$ and $ccss\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{cc}104.8&-27.2\\ -27.2&107.2\end{array}\right)$ $\left(\begin{array}[]{c}133.2\\ 78.8\end{array}\right)$ [$\left[\begin{array}[]{cc}\{-0.69,0.72\}\\ \{-0.72,-0.69\}\end{array}\right]$ $\left(\begin{array}[]{c}4697.0\\ 4642.6\end{array}\right)$ $\left(\begin{array}[]{c}4774.7\\ 4720.3\end{array}\right)$ $\left(\begin{array}[]{c}4758.3\\ 4703.9\end{array}\right)$ $\left(\begin{array}[]{c}5028.0\\ 4973.6\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccccccc}-107.0&38.6&114.7&-29.4&-49.0&125.7&0.0\\ 38.6&-31.1&-29.4&21.9&125.7&-19.6&59.9\\ 114.7&-29.4&7.1&-12.3&20.5&-23.0&-0.8\\ -29.4&21.9&-12.3&57.4&-23.0&8.2&-6.4\\ -49.0&125.7&20.5&-23.0&48.0&0.0&-16.6\\ 125.7&-19.6&-23.0&8.2&0.0&-70.4&-0.8\\ 0.0&59.9&-0.8&-6.4&-16.6&-0.8&26.3\end{array}\right)$ $\left(\begin{array}[]{c}149.7\\ 106.0\\ 56.0\\ 40.8\\ -10.6\\ -121.9\\ -289.8\end{array}\right)$ [$\left[\begin{array}[]{ccccccc}\{0.19,-0.58,0.14,-0.01,-0.74,0.15,-0.18\}\\ \{-0.52,-0.12,-0.61,0.48,-0.20,-0.26,-0.08\}\\ \{0.29,0.27,0.13,0.70,-0.18,0.26,0.49\}\\ \{-0.09,0.04,-0.29,-0.50,-0.30,-0.09,0.75\}\\ \{0.24,0.13,-0.62,-0.14,0.03,0.69,-0.21\}\\ \{0.25,0.68,-0.07,-0.10,-0.46,-0.37,-0.33\}\\ \{-0.69,0.33,0.35,-0.10,-0.25,0.46,-0.07\}\end{array}\right]$ $\left(\begin{array}[]{c}4713.5\\ 4669.8\\ 4619.8\\ 4604.6\\ 4553.2\\ 4441.9\\ 4274.0\end{array}\right)$ $\left(\begin{array}[]{c}4791.2\\ 4747.5\\ 4697.5\\ 4682.3\\ 4630.9\\ 4519.6\\ 4351.7\end{array}\right)$ $\left(\begin{array}[]{c}4774.8\\ 4731.1\\ 4681.1\\ 4665.9\\ 4614.5\\ 4503.2\\ 4335.3\end{array}\right)$ $\left(\begin{array}[]{c}5044.5\\ 5000.8\\ 4950.8\\ 4935.6\\ 4884.2\\ 4772.9\\ 4605.0\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccccccc}-62.9&1.7&28.8&145.1&-37.2&31.0&-79.5&0.0\\ 1.7&18.7&1.5&-37.2&27.7&-79.5&12.4&-37.9\\ 28.8&1.5&-84.0&0.0&0.0&0.0&65.6&10.7\\ 145.1&-37.2&0.0&-120.0&27.2&-45.3&145.6&-0.8\\ -37.2&27.7&0.0&27.2&-25.6&145.6&-18.1&-86.8\\ 31.0&-79.5&0.0&-45.3&145.6&74.8&0.0&-16.6\\ -79.5&12.4&65.6&145.6&-18.1&0.0&-124.0&-0.8\\ 0.0&-37.9&10.7&-0.8&-86.8&-16.6&-0.8&82.5\end{array}\right)$ $\left(\begin{array}[]{c}216.9\\ 124.3\\ 76.8\\ -3.0\\ -22.5\\ -91.7\\ -154.3\\ -387.0\end{array}\right)$ [$\left[\begin{array}[]{cccccccc}\{0.02,0.12,0.03,0.07,-0.56,-0.70,0.07,0.42\}\\ \{-0.17,0.60,-0.05,-0.10,0.12,-0.40,0.01,-0.65\}\\ \{0.53,-0.06,0.22,0.68,0.04,-0.07,0.35,-0.26\}\\ \{0.59,-0.08,-0.35,-0.08,-0.19,-0.05,-0.66,-0.20\}\\ \{0.32,0.73,0.14,0.01,0.24,0.22,-0.10,0.48\}\\ \{-0.14,0.03,-0.83,0.36,0.31,-0.11,0.12,0.20\}\\ \{0.01,0.27,-0.31,-0.03,-0.66,0.50,0.35,-0.15\}\\ \{-0.46,0.12,0.16,0.61,-0.21,0.18,-0.54,-0.03\}\end{array}\right]$ $\left(\begin{array}[]{c}4780.7\\ 4688.1\\ 4640.6\\ 4560.8\\ 4541.3\\ 4472.1\\ 4409.5\\ 4176.8\end{array}\right)$ $\left(\begin{array}[]{c}4858.4\\ 4765.8\\ 4718.3\\ 4638.5\\ 4619.0\\ 4549.8\\ 4487.2\\ 4254.5\end{array}\right)$ $\left(\begin{array}[]{c}4842.0\\ 4749.4\\ 4701.9\\ 4622.1\\ 4602.6\\ 4533.4\\ 4470.8\\ 4238.1\end{array}\right)$ $\left(\begin{array}[]{c}5111.7\\ 5019.1\\ 4971.6\\ 4891.8\\ 4872.3\\ 4803.1\\ 4740.5\\ 4507.8\end{array}\right)$

### v2

@@ III.4 Spectra and decay widths of $ccss\bar{n}$ and $ccss\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{cc}104.8&-27.2\\ -27.2&107.2\end{array}\right)$ $\left(\begin{array}[]{c}133.2\\ 78.8\end{array}\right)$ $\left[\begin{array}[]{cc}\{-0.69,0.72\}\\ \{-0.72,-0.69\}\end{array}\right]$ $\left(\begin{array}[]{c}4697.0\\ 4642.6\end{array}\right)$ $\left(\begin{array}[]{c}4774.7\\ 4720.3\end{array}\right)$ $\left(\begin{array}[]{c}4758.3\\ 4703.9\end{array}\right)$ $\left(\begin{array}[]{c}5028.0\\ 4973.6\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccccccc}-107.0&38.6&114.7&-29.4&-49.0&125.7&0.0\\ 38.6&-31.1&-29.4&21.9&125.7&-19.6&59.9\\ 114.7&-29.4&7.1&-12.3&20.5&-23.0&-0.8\\ -29.4&21.9&-12.3&57.4&-23.0&8.2&-6.4\\ -49.0&125.7&20.5&-23.0&48.0&0.0&-16.6\\ 125.7&-19.6&-23.0&8.2&0.0&-70.4&-0.8\\ 0.0&59.9&-0.8&-6.4&-16.6&-0.8&26.3\end{array}\right)$ $\left(\begin{array}[]{c}149.7\\ 106.0\\ 56.0\\ 40.8\\ -10.6\\ -121.9\\ -289.8\end{array}\right)$ $\left[\begin{array}[]{ccccccc}\{0.19,-0.58,0.14,-0.01,-0.74,0.15,-0.18\}\\ \{-0.52,-0.12,-0.61,0.48,-0.20,-0.26,-0.08\}\\ \{0.29,0.27,0.13,0.70,-0.18,0.26,0.49\}\\ \{-0.09,0.04,-0.29,-0.50,-0.30,-0.09,0.75\}\\ \{0.24,0.13,-0.62,-0.14,0.03,0.69,-0.21\}\\ \{0.25,0.68,-0.07,-0.10,-0.46,-0.37,-0.33\}\\ \{-0.69,0.33,0.35,-0.10,-0.25,0.46,-0.07\}\end{array}\right]$ $\left(\begin{array}[]{c}4713.5\\ 4669.8\\ 4619.8\\ 4604.6\\ 4553.2\\ 4441.9\\ 4274.0\end{array}\right)$ $\left(\begin{array}[]{c}4791.2\\ 4747.5\\ 4697.5\\ 4682.3\\ 4630.9\\ 4519.6\\ 4351.7\end{array}\right)$ $\left(\begin{array}[]{c}4774.8\\ 4731.1\\ 4681.1\\ 4665.9\\ 4614.5\\ 4503.2\\ 4335.3\end{array}\right)$ $\left(\begin{array}[]{c}5044.5\\ 5000.8\\ 4950.8\\ 4935.6\\ 4884.2\\ 4772.9\\ 4605.0\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccccccc}-62.9&1.7&28.8&145.1&-37.2&31.0&-79.5&0.0\\ 1.7&18.7&1.5&-37.2&27.7&-79.5&12.4&-37.9\\ 28.8&1.5&-84.0&0.0&0.0&0.0&65.6&10.7\\ 145.1&-37.2&0.0&-120.0&27.2&-45.3&145.6&-0.8\\ -37.2&27.7&0.0&27.2&-25.6&145.6&-18.1&-86.8\\ 31.0&-79.5&0.0&-45.3&145.6&74.8&0.0&-16.6\\ -79.5&12.4&65.6&145.6&-18.1&0.0&-124.0&-0.8\\ 0.0&-37.9&10.7&-0.8&-86.8&-16.6&-0.8&82.5\end{array}\right)$ $\left(\begin{array}[]{c}216.9\\ 124.3\\ 76.8\\ -3.0\\ -22.5\\ -91.7\\ -154.3\\ -387.0\end{array}\right)$ $\left[\begin{array}[]{cccccccc}\{0.02,0.12,0.03,0.07,-0.56,-0.70,0.07,0.42\}\\ \{-0.17,0.60,-0.05,-0.10,0.12,-0.40,0.01,-0.65\}\\ \{0.53,-0.06,0.22,0.68,0.04,-0.07,0.35,-0.26\}\\ \{0.59,-0.08,-0.35,-0.08,-0.19,-0.05,-0.66,-0.20\}\\ \{0.32,0.73,0.14,0.01,0.24,0.22,-0.10,0.48\}\\ \{-0.14,0.03,-0.83,0.36,0.31,-0.11,0.12,0.20\}\\ \{0.01,0.27,-0.31,-0.03,-0.66,0.50,0.35,-0.15\}\\ \{-0.46,0.12,0.16,0.61,-0.21,0.18,-0.54,-0.03\}\end{array}\right]$ $\left(\begin{array}[]{c}4780.7\\ 4688.1\\ 4640.6\\ 4560.8\\ 4541.3\\ 4472.1\\ 4409.5\\ 4176.8\end{array}\right)$ $\left(\begin{array}[]{c}4858.4\\ 4765.8\\ 4718.3\\ 4638.5\\ 4619.0\\ 4549.8\\ 4487.2\\ 4254.5\end{array}\right)$ $\left(\begin{array}[]{c}4842.0\\ 4749.4\\ 4701.9\\ 4622.1\\ 4602.6\\ 4533.4\\ 4470.8\\ 4238.1\end{array}\right)$ $\left(\begin{array}[]{c}5111.7\\ 5019.1\\ 4971.6\\ 4891.8\\ 4872.3\\ 4803.1\\ 4740.5\\ 4507.8\end{array}\right)$
@@ V ACKNOWLEDGMENTS
Data Availability Statement No Data were associated with the manuscript.